If 26.3 mL of a fruit juice requires 13.7 mL of 0.0100 M I2 solution for reaction, what is the molarity of the ascorbic acid in the fruit juice?
Answer #1

V1 = 26.3 ml

M2 = ?

V2 = 13.7 ml

M = 0.01 I2

Assume 1 mol of I2 reacts with 1 mol of fruit juice

Ascorbic acid + I2 <--> 2HI + ascorbate

then

mol of acid = mol of I2

mol of I2 = MV = 13.7*0.01 = 0.137 mmol of I2

now,

since ratio is 1:1

0.137 mmol of acid was there

[Acids] = mmol/V = 0.137/26.3 = 0.005209

[Ascorbic Acid] = 0.005209 M